40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Oct 29 14:39:37 2021, Anchal, Summary, BHD, 1Y1 cleared. IY3 ready for C1SUS2 I/O and FE. RemovedStuff.jpeg1Y1.jpeg1Y3.jpeg
    Reply  Tue Nov 2 16:42:00 2021, Paco, Summary, BHD, 1Y1 rack work 
       Reply  Tue Nov 2 18:54:41 2021, Anchal, Paco, Update, BHD, Optical fibres laid for BHD upgrade 
       Reply  Thu Nov 4 15:03:43 2021, Koji, Summary, BHD, 1Y1 rack work 
          Reply  Mon Nov 8 10:13:52 2021, Paco, Summary, BHD, 1Y1 rack work; Sorensens removed 
             Reply  Mon Nov 8 13:13:00 2021, Koji, Summary, BHD, 1Y1 rack work; Sorensens removed 40m_BHD.png
                Reply  Mon Nov 8 15:29:05 2021, Paco, Summary, BHD, 1Y1 rack work; New power for cameras 
                   Reply  Mon Nov 8 18:42:38 2021, Koji, Summary, BHD, Rack Layout / Power Strips rack_plan.pdf
                      Reply  Tue Nov 9 19:02:47 2021, Anchal, Summary, BHD, 1Y0 Populated and 1Y1,1Y0 powered 
Message ID: 16448     Entry time: Thu Nov 4 15:03:43 2021     In reply to: 16444     Reply to this: 16453
Author: Koji 
Type: Summary 
Category: BHD 
Subject: 1Y1 rack work 

I have visited the binder file for the 40m wiring file in the control room.
The 12V power supply on 1Y1 is for the CCD cameras. So we still want to keep the 12V 0.8A power and the side connections for these. It is not necessary to be Sorensen. Can we replace it with an AC-DC adapter with +12V/1A for example? BTW, the video matrix and quads are AC-powered.

The mysterious thick cables and cross-connects (green wires) on the side panel (labeled AP1/AP2/SP/IMCREFL) are for "EO shutters". It was meant for the protection of the PDs from bright beams.
I don't think they have been used. And we don't need them.

ELOG V3.1.3-